OpenStudy (anonymous):

Solve cos +√ 2=-cos x for x over the interval [0, 2π]. Please help.

OpenStudy (anonymous):

\[ \cos +√ 2=-\cos x\]\[√ 2=-2\cos x\]\[-√ 2~/~2=\cos x\]\[x=-45º~~is~~one~~solution\]there can be more.

OpenStudy (anonymous):

I think....

OpenStudy (anonymous):

The answers are either A)3π\4 B)x=3π/4+2nπ C)3π\4 and 5π/4 D)x=π\4 and 7π\4